# Offline Fictitious Self-Play for Competitive Games

### v1

@@ 4 Offline Fictitious Self-Play
In this section, we introduce the Offline Self-Play (Off-SP) learning framework and give an implementation as Offline Fictitious Self-Play (Off-FSP) algorithm to find an $\epsilon$-Nash equilibrium with a fixed dataset. In the online scenario, self-play is a process in which a strategy iteratively maximizes cumulative rewards against itself, and fictitious self-play plays against the average of its past strategies. Corresponding to the concept of the online setting, Off-SP is a learning framework that enables strategy play against changing opponents with a fixed dataset. Off-FSP implement fictitious self-play on this learning framework with average strategies as the changing opponents and learn to find NE.
@@
Initially, we describe the offline datasets and make strong assumptions to simplify the problem. Based on the original FSP, as described in Algorithm 1, we derive the offline fictitious self-play by replacing three essential functions. In Section 4.2, we replace GenerateData by simulating play against different opponents with weighted datasets and replace LearnBestResponse with single-agent offline RL algorithm. The first two functions introduce Off-SP. In Section 4.3, we replace UpdateAverageStrategy by computing the average strategy based on samples and derive Off-FSP. Lastly, we discuss how offline RL helps Off-FSP to reduce the exploitability under real-world datasets.
4.1 Assumptions of Offline Datasets
Trajectory and Dataset.
The trajectory of extensive-form games is a sequence of information states, actions, and player’s id. The trajectory also involves states and actions of chance $c$. The trajectory $\tau_{E}=\{s_{1},a_{1},p_{1},s_{2},\dots,s_{T},a_{T},p_{T}\mid p_{t}=P(s_{t})\},$
We can assume the next information state for a prefix of trajectory and the payoff for a whole trajectory are definite. In other words, if we know the information sets of all players and chance, the game is deterministic.
@@ 4.1 Assumptions of Offline Datasets
In player $i$’s perspective, the extensive-form game can be modeled as an MDP $\mathcal{M}$ given a fixed opponent $\pi^{-i}$ (Silver & Veness, 2010; Greenwald et al., 2013). In MDP $\mathcal{M}$, the dynamic function $\mathcal{T}(s^{i}_{t+1}\mid s^{i}_{t},a^{i}_{t})$ models the dynamics of opponents and the chance player. The observed MDP is non-deterministic because of imperfect information in a single player’s information state.
@@
To learn best-response with single-agent RL, we project the trajectory $\tau_{E}$ into player $i$’s perspective with a projection function $\tau^{i}=\mathcal{F}^{i}(\tau_{E})$. The projection $\mathcal{F}^{i}$ filters out the states corresponding to player $i$ while relabeling the time indices.
$\begin{split}\tau^{i}&=\mathcal{F}^{i}(\tau_{E})=\{s_{t},a_{t}\mid\forall p_{t}\in\tau_{E},p_{t}=i\}\\ &=\{s_{1}^{i},a_{1}^{i},s_{2}^{i},\dots,s_{T^{\prime}}^{i},a_{T^{\prime}}^{i}\},\end{split}$
@@
where $T^{\prime}$ is the length of the player $i$’s trajectory $\tau^{i}_{\mathcal{M}}$. The subscripts of $s_{t}$ and $s_{t}^{i}$ are different and represent the time indices in the corresponding trajectories. For state $s^{i}_{t}\in\tau^{i}_{\mathcal{M}}$, we use function $I(s^{i}_{t})$ to denote the subscript of state in $\tau_{E}$. We denote opponent $j$’s latest state at player $i$’s state $s_{t}^{i}$ with function $\tau_{<}^{j}(s_{t}^{i})$.
@@
In order to offer a clear explanation of the stated components, we present an example trajectory in Figure 2.
Figure 2: An example of extensive-form games and datasets. (Left) An example trajectory in extensive form game; (Right) Illustration of trajectories and dataset in extensive-form game and single-agent perspective. The orange area represents the player function. Blue parts represent the trajectories. The purple area represents the projection relationship under function $\mathcal{F}^{1}$. The green parts explain how function $\tau^{2}_{<}$ works.
Based on the description of trajectory in both extensive-form games and single-agent perspective, the datasets consist of multiple corresponding trajectories. For extensive-form game, the dataset is $\mathcal{D}_{E}=\{\tau_{E}\}$. The dataset for player $i$ is $\mathcal{D}^{i}=\{\tau^{i}_{\mathcal{M}}=\mathcal{F}^{i}(\tau_{E})\mid\tau_{E}\in\mathcal{D}_{E}\}$. The $i$-th player’s dataset, derived from dataset $\mathcal{D}_{E}$, is denoted as $\mathcal{D}^{i}=\mathcal{F}^{i}(\mathcal{D}_{E})$. For briefly, we ignore the reward transformation in trajectories.
Assumptions for Offline Learning.
To introduce our method, we first assume the offline dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset. This assumption, crucial for the theoretical derivation of our methods, is not a prerequisite for practical application. The adaptability of our method to real-world datasets, unconstrained by these assumptions, is discussed in Section 4.4. The definitions of these two concepts are provided as follows.
@@
[…]
@@
For a dataset of extensive-form game $\mathcal{D}_{E}$ sampled following a behavior strategy $\pi_{b}$, dataset $\mathcal{D}_{E}$ is an real-equivalence dataset if $Pr(\tau_{E})=\frac{\sum_{\tau\in\mathcal{D}_{E}}\mathbbm{1}[\tau=\tau_{E}]}{|\mathcal{D}_{E}|},\forall\tau_{E}\in\mathcal{D}_{E},$ where $Pr(\tau_{E})$ is the probability of sampling trajectory $\tau_{E}$ with $\pi_{b}$ online.
When a dataset is a real-equivalence dataset , sampling trajectories from the offline dataset $\mathcal{D}_{E}$ is equivalent to sampling from the extensive-form game with strategy $\pi_{b}$ online and the right-hand side of the equation is the probability density of $\tau_{E}$ in $\mathcal{D}_{E}$, denoted by $\mathcal{D}_{E}(\tau_{E})$.
4.2 Learning Open-ended Best-Response
Self-play and fictitious self-play iteratively interact with the game to generate data and utilize the data to learn best responses. In $k$-th iteration, the opponent for player $i$ is changed to $\pi^{-i}_{k-1}$, therefore, in order to learn the best response $\beta^{i}_{k}=\text{BR}(\pi^{-i}_{k-1})$, player $i$ must interact with $\pi^{-i}_{k-1}$ to generate new data $\mathcal{D}^{i}_{k}$. The process, from the perspective of player $i$, is equivalent to interacting with a new MDP $\mathcal{M}(\pi^{-i}_{k})$ and maximizing returns with RL to learn the corresponding best response (Greenwald et al., 2013).
In offline settings, we are limited to using a fixed dataset sampled by behavioural policy $\pi_{b}$, in which RL methods can only obtain the best response $\text{BR}(\pi_{b}^{-i})$ for player $i$. In order to execute the fictitious self-play, we should generate player $i$ datasets under different MDP $\mathcal{M}(\pi^{-i}_{k})$. With importance sampling (Nachum et al., 2019), we can emulate sampling from another dataset $\mathcal{D}^{i}_{w}$ with weighting $w(d_{t})=\frac{\mathcal{D}^{i}_{w}(d_{t})}{\mathcal{D}^{i}(d_{t})}$ (Hong et al., 2023) on the original dataset $\mathcal{D}^{i}$, where $\mathcal{D}^{i}_{w}(d_{t})$ and $\mathcal{D}^{i}(d_{t})$ denote the probability density of tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$ in according datasets. This formulation gives the following equivalence:
@@ 4.2 Learning Open-ended Best-Response
where $L(d_{t};\theta)$ is the loss function of an off-policy RL method, and $\theta$ is the parameter of RL policy to be optimized. With the assumption that dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset, sampling from it with importance sampling is equivalent to sampling from the online game with importance sampling.
@@
For player $i$, the weight of transferring opponent from $\pi^{-i}_{b}$ to $\pi^{-i}$ is:
@@
In order to simplify the phrase, we denote the strategies of $n-1$ opponents as an ensemble strategy $\pi^{-i}$ and refer to them as player $-i$. The proof of Theorem 4.3 is provided in Appendix A. Given an offline dataset $\mathcal{D}_{E}$, the behavior strategy $\pi_{b}$ can be approximated by an empirical behavioural policy $\tilde{\pi}_{b}$. In extensive-form games, we can estimate $\tilde{\pi}_{b}$ by counting or supervised learning policy.
@@
When considering a particular opponent, the weight $w(d_{t})$ assigned to a sample $d_{t}$ may shift to zero or a large value. As a result, when training RL with batched samples, the RL loss suffers from a large variance (Munos et al., 2016), leading to unstable training. To address this issue, we employ $\frac{w(d_{t})}{\sum_{d\in\mathcal{D}}w(d)}$ as the sampling probability rather than multiply the loss function by $w(d_{t})$ as shown in Equation 2. In this way, re-sampling from fully covered real-equivalence datasets is identical to sample data in online games. With batches of data sampled from the dataset, we can apply an offline RL to learn best responses.
@@
Figure 0(b) and Figure 3 illustrates these two functions and Off-SP. In the function GenerateDataset, we first calculate $w(d)$ for all the samples $d_{t}$ and get a re-weighted dataset $\mathcal{D}_{k}^{i}$ for each player $i$. In the function LearnBestResponse, offline RL algorithms repeatedly sample a batch of data and optimize the learned policy for $M$ times. These two functions derive the Off-SP. At each step $t$, Off-SP first generates dataset play against the current opponents and updates strategies towards the best response of the opponents.
Figure 3: Offline Self-Play learning framework.
4.3 Aggregate Average Strategy in Samples
Before diving into Off-FSP, it is essential to review the mechanics of fictitious play in an online scenario. In norm-form games, fictitious play updates the new average strategy $\pi_{k+1}$ following Equation 1 after learning the best responses. To extend FP on extensive-form games, FSP gathers the trajectories of $\beta^{i}_{k}$ during each iteration and employs supervised learning to approximate $\pi^{i}_{k+1}$ for each player. However, Off-FSP does not require the current average strategy $\pi_{k}$ to sample new data, but only requires probability $\pi_{k}(s,a)$ for all state-action pairs in a fixed dataset. Therefore, an interactable average strategy $\pi_{k}$ is not necessary. We can just maintain the probability $\pi_{k}(s,a)$ in the datasets for the weighting technique. Traverse along the trajectory $\tau_{i}$, Off-FSP computes the probability $\pi^{i}_{k}(s,a)$ for player $i$ following Equation 1. Figure 0(b) shows one step of the function UpdateAverageStrategy. To facilitate the calculation, we save the probability into the current dataset $\mathcal{D}^{i}_{k}$, changing the sample into $d^{\prime}=(s,a,\pi^{i}_{k}(s,a))$.
Although the weighting technique does not require a real instance of average behavioural strategy, Off-FSP requires an interactable strategy as output. One way is utilizing supervised learning to imitate the behavioural strategy with the weighting technique. When the dataset is not a fully covered dataset, this restricts strategies to exclusively choosing actions that are present in the dataset. In our experiments, we opted for an alternative method that involves aggregating all policies within a collection $\Pi$. Before the beginning of each game, one of the strategies in $\Pi$ is chosen with a specific probability and the payoff of the aggregation is the expectative payoff over all possible strategies. This approach maintains the option of picking actions that are out of the distribution but necessitates a larger storage capacity for storing all strategies. Algorithm 2 and Figure 0(a) present the pipeline of Offline Fictitious Self-play.

### v2

@@ 4 Offline Fictitious Self-Play
In this section, we introduce the Offline Self-Play (Off-SP) learning framework and give an implementation as Offline Fictitious Self-Play (Off-FSP) algorithm to minimize NashConv, the distance to NE, with a fixed dataset. Off-SP learns policies iteratively, maximizing cumulative rewards against changing opponents without interactable environments. Off-FSP adopts the fictitious self-play on Off-SP, where the policy plays against the average of past opponents.
@@
Initially, we describe the offline datasets and make assumptions to simplify the problem. Based on the original FSP, as described in Appendix˜A, we derive the offline fictitious self-play with modifications on three essential functions, GenerateData, LearnBestResponse, and UpdateAveragePolicy. In Off-SP, GenerateData simulates play against different opponents with weighted datasets, and LearnBestResponse optimizes the policy with an existing single-agent offline RL algorithm. Offline RL algorithm, such as CQL, ensures the performance of BR and reduction of NashConv even on partially covered datasets. In Section˜4.3, we introduce UpdateAveragePolicy by computing the average policy on samples and derive Off-FSP.
4.1 Problem Formulation
@@
The trajectory of extensive-form games is a sequence of information states, actions, and rewards. The trajectory is $\tau_{E}=\{s_{1},a_{1},r_{1},\dots,s_{T},a_{T},r_{T}\}$. In player $i$’s perspective, the extensive-form game can be modeled as an MDP $\mathcal{M}$ given a fixed opponent $\pi^{-i}$ (Silver and Veness 2010; Greenwald et al. 2013). In MDP $\mathcal{M}$, the dynamic function $\mathcal{T}(s^{i}_{t+1}\mid s^{i}_{t},a^{i}_{t})$ models the dynamics of opponents.
@@
The goal of Off-FSP is to iteratively learn the best-response policy $\pi^{i}$ against changing opponents and minimize NashConv. To learn best-response with single-agent RL, we project the trajectory $\tau_{E}$ into player $i$’s perspective with a projection function $\tau^{i}=\mathcal{F}^{i}(\tau_{E})$. The projection $\mathcal{F}^{i}$ filters out the states corresponding to player $i$ while relabeling the time indices.
$\begin{split}\tau^{i}=\mathcal{F}^{i}(\tau_{E})&=\{s_{t},a_{t},r_{t}\mid\forall s_{t}\in\tau_{E},\mathcal{P}(s_{t})=i\}\\ &=\{s_{1}^{i},a_{1}^{i},r_{1}^{i},s_{2}^{i},\dots,s_{T^{\prime}}^{i},a_{T^{\prime}}^{i},r_{T^{\prime}}^{i}\},\end{split}$
where $T^{\prime}$ is the length of the player $i$’s trajectory $\tau^{i}_{\mathcal{M}}$.
@@
The subscripts of $s_{t}$ and $s_{t}^{i}$ are different and represent the time indices in the corresponding trajectories. For state $s^{i}_{t}\in\tau^{i}_{\mathcal{M}}$, we use function $I(s^{i}_{t})$ to denote the subscript of state in $\tau_{E}$, and $\tau_{<}^{j}(s_{t}^{i})$ denotes opponent $j$’s latest state at $s_{t}^{i}$.
@@
In order to offer a clear explanation of the stated components, we present an example trajectory in Figure 5. Based on the description of trajectory in both extensive-form games and single-agent perspective, the datasets consist of multiple corresponding trajectories. For extensive-form game, the dataset is $\mathcal{D}_{E}=\{\tau_{E}\}$. The dataset for player $i$ is $\mathcal{D}^{i}=\{\tau^{i}_{\mathcal{M}}=\mathcal{F}^{i}(\tau_{E})\mid\tau_{E}\in\mathcal{D}_{E}\}$. The $i$-th player’s dataset, derived from dataset $\mathcal{D}_{E}$, is denoted as $\mathcal{D}^{i}=\mathcal{F}^{i}(\mathcal{D}_{E})$.
Figure 5: An illustration example of trajectory $\tau^{1}$ and $\tau_{E}$. Purple parts represent Player 1. The yellow arrows imply the projection relationship under function $\mathcal{F}^{1}$. Green part indicates $\tau_{<}^{1}$.
To introduce the theoretical foundation of our method, we begin by defining an idealized dataset condition that facilitates analysis. These assumptions are not required for practical application and are used only to isolate and clarify the core mechanisms of our approach. In later sections, we show how the method generalizes to realistic datasets with partial coverage by integrating standard offline RL algorithms.
@@
$\mathcal{S}\text{~and~}\mathcal{A}$ represent the joint sets of information states and actions for all players $\mathcal{N}$. A dataset $\mathcal{D}_{E}$ in extensive-form games is a fully covered dataset if $\forall s\in\mathcal{S},a\in\mathcal{A}(s)\Rightarrow(s,a)\in\mathcal{D}_{E}$.
@@
For a dataset of extensive-form game $\mathcal{D}_{E}$ sampled following a policy $\pi_{b}$, dataset $\mathcal{D}_{E}$ is an real-equivalence dataset if $Pr(\tau_{E})=\frac{\sum_{\tau\in\mathcal{D}_{E}}\mathds{1}[\tau=\tau_{E}]}{|\mathcal{D}_{E}|},\forall\tau_{E}\in\mathcal{D}_{E},$ where $Pr(\tau_{E})$ is the probability of sampling trajectory $\tau_{E}$ with $\pi_{b}$.
When a dataset is a real-equivalence dataset , sampling trajectories from the offline dataset $\mathcal{D}_{E}$ is equivalent to sampling from the extensive-form game with policy $\pi_{b}$ online and the right-hand side of the equation is the probability density of $\tau_{E}$ in $\mathcal{D}_{E}$, denoted by $\mathcal{D}_{E}(\tau_{E})$.
4.2 Offline Self-Play
Self-play iteratively interacts with the game to generate data and utilizes the data to learn the best responses. In $k$-th iteration, the opponent for player $i$ is changed to $\pi^{-i}_{k-1}$, therefore, in order to learn the best response $\beta^{i}_{k}=\text{BR}(\pi^{-i}_{k-1})$, player $i$ must interact with $\pi^{-i}_{k-1}$ to generate new data $\mathcal{D}^{i}_{k}$. The process of learning BR, from the perspective of player $i$, is equivalent to interacting with a new MDP $\mathcal{M}(\pi^{-i}_{k})$ and maximizing returns with RL (Greenwald et al. 2013).
In offline settings, we are limited to using a fixed dataset sampled by behavioural policy $\pi_{b}$, in which RL methods can only obtain the best response $\text{BR}(\pi_{b}^{-i})$ for player $i$. In order to execute the fictitious self-play, we generate player $i$’s datasets under different MDP $\mathcal{M}(\pi^{-i}_{k})$. With importance sampling (Nachum et al. 2019), we can emulate sampling from another dataset $\mathcal{D}^{i}_{w}$ with weighting $w(d_{t})=\frac{\mathcal{D}^{i}_{w}(d_{t})}{\mathcal{D}^{i}(d_{t})}$ (Hong et al. 2023) on the original dataset $\mathcal{D}^{i}$, where $\mathcal{D}^{i}_{w}(d_{t})$ and $\mathcal{D}^{i}(d_{t})$ denote the probability density of tuple $d_{t}=(s^{i}_{t},a^{i}_{t},s^{i}_{t+1})$. This formulation gives the following equivalence:
@@ 4.2 Offline Self-Play
where $L(d_{t};\theta)$ is the loss function of an off-policy RL method, and $\theta$ is the parameter of the RL policy to be optimized. With the assumption that dataset $\mathcal{D}_{E}$ is both a fully covered dataset and a real-equivalence dataset, sampling from it with importance sampling is equivalent to sampling from the online game with importance sampling.
@@
For player $i$, the weight of transferring the opponent from $\pi^{-i}_{b}$ to $\pi^{-i}$ is:
@@
The proof of Theorem 4.3 is provided in Appendix C. Given an offline dataset $\mathcal{D}_{E}$, the policy $\pi_{b}$ can be approximated by an empirical behavioural policy $\tilde{\pi}_{b}$. We can estimate $\tilde{\pi}_{b}$ by counting or supervised learning policy.
@@
In the learning process, the weight $w(d_{t})$ assigned to a sample $d_{t}$ may shift to zero or a large value. As a result, the RL loss suffers from a large variance (Munos et al. 2016), leading to unstable training. To address this issue, we employ $\frac{w(d_{t})}{\sum_{d\in\mathcal{D}}w(d)}$ as the sampling probability rather than multiplying the loss function by $w(d_{t})$ as shown in Equation 2. In this way, re-sampling from fully covered real-equivalence datasets is identical to sampling data in online games. With batches of data sampled from the dataset, we can apply an offline RL to learn BR.
@@
Figures˜4(c) and 4(a) illustrates these two functions and Off-SP. In the function GenerateData, we first calculate $w(d)$ for all the samples $d_{t}$ and get a re-weighted dataset $\mathcal{D}_{k}^{i}$ for each player $i$. In the function LearnBestResponse, offline RL algorithms repeatedly sample a batch of data and optimize the learned policy for $M$ times. We use CQL (Kumar et al. 2020) as the default offline RL algorithm in our implementation, i.e., $\beta^{i}_{k}=\text{CQL}(\mathcal{D}_{k}^{i})$. These two functions derive the Off-SP. At each step $t$, Off-SP first generates a dataset playing against the current opponents and updates policies towards the best response of the opponents.
Figure 6: Results on Extensive-Form Games. (Top) NashConv on Mix Datasets; (Bottom) NashConv on Population Datasets.
Figure 7: Coverage ratio of datasets.
4.3 Aggregate Average Policy in Samples
Similar to FSP, Off-FSP play against the average policy $\pi_{k+1}$ following  Equation˜1. Although an interactable average policy $\pi_{k}$ is not necessary in the offline setting, we can just maintain the probability $\pi_{k}(s,a)$ in the datasets for the weighting technique. Traverse along the trajectory $\tau_{i}$, Off-FSP computes the probability $\pi^{i}_{k}(s,a)$ for player $i$ following Equation˜1. Figure˜4(c) shows one step of the function UpdateAveragePolicy. To facilitate the calculation, we save the probability into the current dataset $\mathcal{D}^{i}_{k}$, changing the sample into $d^{\prime}=(s,a,\pi^{i}_{k}(s,a))$. In short, we have:
$\pi^{i}_{k}(s,a)\leftarrow\frac{k-1}{k}x^{i}_{\pi^{i}_{k-1}}(s)\pi^{i}_{k-1}(s,a)+\frac{1}{k}x^{i}_{\beta^{i}_{k}}(s)\beta^{i}_{k}(s,a).$
In the evaluation phase, we aggregate all policies within a collection $\Pi$. Before the beginning of each game, one of the policies in $\Pi$ is chosen with a specific probability, and the payoff of the aggregation is the expected payoff over all possible policies. To keep the storage efficient, we only keep policies at fixed interval steps for the real-world application in Section˜5.2. Algorithm 2 in Appendix˜A and Figure˜4(b) presents the pipeline of Offline Fictitious Self-play.
Memory and Computational Complexity. Compared with offline RL, Off-FSP only assigns additional weights for samples in the dataset, so the memory complexity is still $O(|\mathcal{D}_{E}|)$ in training. The training time of Off-FSP can be noted as $T=T_{reweight}+T_{learn}$, where $T_{reweight}$ is the time of re-weighting and $T_{learn}$ is the time of offline RL learning. In empirical experiments of Section˜5, the $T_{reweight}:T_{learn}\approx 1:1$.
5 Experiments
In this section, we design experiments to show the performance of Off-FSP in offline competitive games. Three benchmark extensive-form games are selected for analysis: Leduc Poker, Large Kuhn Poker, and Oshi Zumo, which include two stochastic poker games and one deterministic board game. These environments allow efficient sampling and easy NashConv computation, facilitating a comprehensive assessment of algorithm performance. We compare Off-FSP with state-of-the-art baselines, including OEF (Li et al. 2022) and Behavior Cloning (BC). We further conduct ablations on offline RL components to demonstrate the flexibility of Off-FSP. Finally, the practical applicability of Off-FSP is highlighted through deployment in a real-world human-robot competitive task, showcasing its potential for solving complex, hard-to-simulate problems. Further experimental details are provided in the appendix.
5.1 Extensive-form Games
To further evaluate Off-FSP on complicated extensive-form games, we collected datasets of different quality on Leduc Poker, Large Kuhn Poker, and Oshi Zumo. Following the previous work (Li et al. 2022), the first type of datasets is called mix datasets, which are sampled from a mixture of expert and random policies. Evaluation on random, mixed, and expert datasets is also a common practice in single-agent offline RL (Kumar et al. 2020). We sample 5 different mix datasets for each game, the ratio of sampling from the expert are $0,0.25,0.5,0.75,1$. The expert policy is learned by online PSRO (Lanctot et al. 2017) with 40 iterations. The second type of dataset is called the population dataset, which is sampled from the population of online PSRO. $5$ population datasets are uniformly sampled from all population policies of $1,10,20,30,40$ iterations of online PSRO. This type of dataset is designed to simulate the datasets sampled from a population of people with different levels of expertise and is similar to the setting in the real world. The population datasets with $1$ iteration and the mix datasets with a ratio of 0 are random datasets. Every dataset comprises 10 000 trajectories, which is far less than the number in the online PSRO. We visualize the coverage of terminal states, i.e., leaf nodes, in Figure˜7. Most of the datasets are partially covered.
Figure˜7 shows the NashConv of Off-FSP and baselines on three extensive-form games. The difficulty of learning from datasets is affected by two factors: the quality of sampling policies and the coverage of datasets. NashConvs of BC show the quality of sampling policies. In most cases, Off-FSP shows the best performance, and the performance of Off-FSP is also robust to both the quality of sampling policies and the coverage of datasets. Both OEF-DCFR and OEF-PSRO, two variants of OEF (Li et al. 2022), fail in most cases. OEF tries to find an offline equilibrium with the model-based paradigm, but it is easy to be misled by O.O.D. actions. OEF mixed its policy with BC, evaluated it in online games multiple times, and used the minimum NashConv as a result. To make a fair comparison, we removed the mixing operation in OEF from our main experiment. The results with mixing are shown in  Appendix˜G. Compared with BC, Off-FSP outperforms in non-expert datasets and achieves comparable performance in expert datasets. In expert datasets, the diversity of samples is limited, and few samples can be exploited to improve the policy. Off-FSP aims to improve the policy with non-expert datasets, which is more practical and easier to scale up in real-world problems.
Figure 8: Results of ablation study on random and mix data of Leduc Poker.
Ablation Studies. In previous experiments, we showed the performance of Off-FSP with CQL as the offline RL algorithm to learn the best response. However, the choice of an offline RL algorithm is also an important factor. Figure˜8 shows learning curve of Off-FSP with different RL algorithms, including CQL, BCQ (Fujimoto, Meger, and Precup 2019), CRR (Wang et al. 2020), and DQN on two datasets of Leduc Poker. Without the constraint of offline RL to avoid O.O.D. actions, Off-FSP with DQN converges to a policy with high exploitability. Corresponding to the design of Off-FSP, it has the potential to combine with any offline RL algorithms. Off-FSP with CQL, BCQ, and CRR show similar performances in these two datasets.
5.2 Real-world Human-Robot Confrontation
Figure 9: Setups of the human-robot competitive game. G1 to G4 in sub-figure (d) are four actions of the human player.
To further evaluate Off-FSP, we design a complex real-world zero-sum game—a human-robot ball defence task, which serves as a simplified version of football defense and attack. Figures˜1 and 9 illustrate the task scenario. This setup involves a human player, making it difficult to simulate and costly to collect data, thus highlighting the significance of Off-FSP. The game consists of two players: a human thrower, who attempts to throw a ball into one of three baskets (goals), and a robotic arm acting as a goalkeeper that tries to block the ball. The human thrower receives a reward of +1+1 if the ball enters a goal and 1-1 otherwise. As a zero-sum game, the robot receives the negative of the human’s reward and uses a black racket to defend the goals. The robot is equipped with a stereo RGBD camera that estimates the ball’s position in real time (see Figure˜9c). Its observation space includes the ball’s position and velocity, as well as the position of the robot’s end effector. The robot policy outputs target end-effector positions at $60$ Hz. These are executed via inverse kinematics and a PD controller. For the human player, we simplify their action space to four discrete targets: the three baskets (G1 to G3) or outside the basket area (G4), as shown in Figure˜9d. The target is solely determined by the ball’s flight trajectory after release and does not change whether the robot blocks it. Therefore, G4 corresponds to imprecise human throws and represents a poor strategy. The setup of this task is shown in Figure˜9.
We collected a mixed dataset comprising $50$ expert trajectories and $150$ low-quality trajectories. Expert data were collected with a human operator teleoperating the robot arm while another expert performed the throws. The average win rate of the teleoperated robot in these expert trajectories was $64\%$. Low-quality trajectories were collected by deploying the robot with a random policy, which moves to uniformly sampled positions within a constrained action space. The human thrower was not an expert and occasionally missed the targets (G4). This protocol makes low-quality data cheaper to collect, requiring only a non-expert human thrower.
Methods G1 G2 G3 Worst Average
BC 0.55 0.35 0.4 0.35 0.43
BC-Expert 0.45 0.5 0.65 0.45 0.53
CQL 0.4 0.85 0.8 0.4 0.68
Off-FSP-CQL 0.8 0.75 0.9 0.75 0.82
Table 1: Winning rate of robot goalkeeper in ball defense game. The colored numbers represent the worst case between goals.
Since one player is human, we evaluate only the robot’s policy across different algorithms. As NashConv is difficult to compute in real-world settings, we adopt the robot’s win rate as the evaluation metric. We compare Off-FSP-CQL with three baselines: BC, BC-Expert, and CQL. Here, BC-Expert denotes BC policy trained only on trajectories with a reward of +1+1. OEF failed in this task, so we removed it from the comparison.
For each goal, a total of 20 throws were performed collectively by 10 human throwers (2 throws per thrower on average). As shown in Table˜1, Off-FSP-CQL outperforms all baselines with consistently robust performance. The dataset shows an imbalanced distribution $P(\text{G1,G2,G3,G4})=(0.29,0.11,0.32,0.28)$, with relatively fewer throws to the middle goal (G2) and a high proportion of off-target throws (G4). Off-FSP learns a more balanced human policy $(0.33,0.22,0.40,0.04)$, leading to improved robustness against diverse human strategies. By focusing only on high-reward data, BC-Expert and CQL overfit to average opponents in the dataset and fail to defend consistently across all goal targets (G1–G3).
6 Conclusion
In this paper, we study offline multi-agent reinforcement learning for competitive games and propose Off-SP and Off-FSP to enable single-agent offline RL algorithms to be applied in this scenario. We find that Off-FSP can approximate NE even with partially covered datasets. Extensive experiments show that all variants of Off-FSP significantly outperform state-of-the-art baselines in different datasets of multiple two-player zero-sum games.
Appendix A Algorithms
Algorithm˜1 describes Fictitious Self-Play (FSP) in the online setting. Algorithm˜2 is the Offline Fictitious Self-Play. Compared with FSP, the function GenerateData and UpdateAveragePolicy only require two additional traverses of the dataset and skip the time of interacting with the environment, so it does not introduce too much computational overhead.
1
2
3Function FictitiousSelfPlay()
4   Initialize policy $\pi_{0}$.;
5
6 for $k=1$ to $K$ do
7
8    $\mathcal{D}_{k}\leftarrow$ GenerateData($\pi_{k-1}$)
9    foreach player $i\in\mathcal{N}$ do
10       $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}_{k},\pi^{-i}_{k-1}$)
11      end foreach
12    $\pi_{k}\leftarrow$ UpdateAveragePolicy($\pi^{i}_{k-1},\beta^{1}_{k},\dots,\beta^{n}_{k}$)
13   end for
14 return average policy $\pi_{K}$;
15
16 end
Algorithm 1 Fictitious Self-Play
1
2
3Function Off-FictitiousSelfPlay($\mathcal{D}_{E}$):
4   Estimate the empirical behavioural policy $\tilde{\pi}_{b}$;
5
6 $\Pi_{i}\leftarrow\{\pi_{b}^{i}\},\quad\pi_{0}\leftarrow\tilde{\pi}_{b}$;
7
8 $\mathcal{D}^{i}_{0}\leftarrow\{(s,a,\pi^{i}_{0}(s,a))|(s,a)\in\mathcal{F}^{i}(\mathcal{D}_{E})\}$;
9
10 for $k=1$ to $K$ do
11    for each player $i\in\mathcal{N}$ do
12       $\mathcal{D}^{i}_{k}\leftarrow$ GenerateData($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$)
13       $\beta^{i}_{k}\leftarrow$ LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$)
14       $\mathcal{D}^{i}_{k}\leftarrow$ UpdateAveragePolicy($\mathcal{D}^{i}_{k-1},\beta_{k}^{i}$)
15        Update collection of BRs $\Pi_{i}\leftarrow\Pi_{i}\cup\{\beta_{k}^{i}\}$;
16
17      end for
18
19   end for
20 return Aggregate($\Pi$)
21
22
231exFunction GenerateData($\mathcal{D}_{E},\mathcal{D}^{-i}_{k-1}$):
24   Get probability of $\pi_{k-1}^{-i}$ from $\mathcal{D}^{-i}_{k-1}$;
25
26  Calculate $w_{\pi_{k-1}^{-i}}$ following Equation 3;
27
28 $\mathcal{D}^{i}_{k}\leftarrow$ WeightData($\mathcal{D}_{E},w_{\pi_{k}^{-i}}$);
29
30 return $\mathcal{D}^{i}_{k}$
31
32
331exFunction LearnBestResponse($\mathcal{D}^{i}_{k},\beta^{i}_{k-1}$):
34   Initialize policy parameters with $\beta^{i}_{k}\leftarrow\beta^{i}_{k-1}$;
35
36  Optimize $\beta^{i}_{k}$ with offline RL and data $\mathcal{D}^{i}_{k}$;
37
38 return $\beta^{i}_{k}$
39
40
411exFunction UpdateAveragePolicy($\mathcal{D}^{i}_{k-1},\beta^{i}_{k}$):
42 for $(s,a,\pi_{k-1}^{i}(s,a))\in\mathcal{D}^{i}_{k-1}$ do
43      Calculate $\pi_{k}^{i}(s,a)$ with $\beta^{i}_{k},\pi^{i}_{k-1}$ following Equation 1, and update dataset $\mathcal{D}^{i}_{k}$;
44
45   end for
46 return $\mathcal{D}^{i}_{k}$
47
48